1- Calculate activity coefficient \gamma of Mg2+ in a 0.0100 M MgCl2 solution.

a - By linear interpolation data in table 7-1

b - By using extended Debye-Huckel equation

media%2Fe1c%2Fe1cd5542-0726-4883-ba03-61
Answer #1

a) From the data given we find,

Ionic strength(micro Molar) Activity coeffcient -log(activity coefficient)
0.001 0.872 0.059483515
0.005 0.755 0.122053048
0.01 0.69 0.161150909
0.05 0.52 0.283996656
0.1 0.45 0.346787486

Plotting log(activity coefficient), log(\gamma \pm) versus ionic strength, we get

0.4 0.35 0.3 0.25 -log(yt)0.063 In(lonic strength) +0.474 0.2 R2 - 0.971 Series1 0.15 Log. (Series1) 0.1 0.05 0 0 0.02 0.04 0

Based on the curve fit we find,

log(\gamma \pm) = 0.063 * Ionic strength + 0.474

At Ionic strength = 0.01M,

log(\gamma \pm) = 0.063*ln(0.01) + 0.474

= 0.183874

Therefore,

\gamma \pm for Mg2+ from interpolation = exp(-0.183874)

\gamma \pm for Mg2+ from interpolation = 0.83204

b) From Debye-Huckel theory,

log(\gamma \pm) = - A*z1*z2*I^0.5

where A = 0.509,

z1 is the charge on the ion 1

z1 is the charge of the ion 2

I is the ionic strength given by

I =0.5* \sumCizi^2

For MgCl2,

I = 0.5* (0.01*2^2+0.02*1^2)

= 0.03

log(\gamma \pm) = - 0.509*2*1*(0.03)^0.5

= - 0.176323

Therefore,

\gamma \pm for Mg2+ from Debye Huckel theory = exp(-0.176323)

\gamma \pm for Mg2+ from Debye Huckel theory = 0.838347